OpenStudy (anonymous):

determine if (1/2, -4) is a solution to 4x-3y=10

jimthompson5910 (jim_thompson5910):

Plug in x = 1/2 and y = -4 Then evaluate/simplify If you get a true equation, then (1/2, -4) is a solution

OpenStudy (campbell_st):

substitute them both to see if the solution is 10 4(1/2) - 3(-4) = ?? if its 10 then the ordered pair is a solution

OpenStudy (anonymous):

how do you get 12

OpenStudy (anonymous):

4*1/2-(3*-4) =(2+12) =14....sry lhs 14...calculation mistake!

OpenStudy (anonymous):

got it thank you

OpenStudy (anonymous):

welcm...:)